OpenStudy (anonymous):

Solve |2x - 5| = 4 A. {x | x = -4.5 or x = 4.5} B. {x | x = 0.5 or x = 4.5} C. {x | 0.5 < x < 4.5}

OpenStudy (anonymous):

\[|2x - 5| = 4\] solve \[2x-5=-4\] and then solve \[2x-5=4\] to get your two solutions

OpenStudy (anonymous):

A ?